A particle of mass m is constrained to move on


the boundary of an ellipse x^2 + y^2 = 2. Identify


the generalize.d coordinates of this system.


1
Expert's answer
2022-05-22T14:37:33-0400

The generalized coordinate is polar angle "\theta":

"x=\sqrt{2}\cos\theta\\
y=\sqrt{2}\sin\theta"

Then

"x^2+y^2=2(\cos^2\theta+\sin^2\theta)=2"